# label=0.,leftmargin=15pt,labelwidth=10pt,labelsep=5pt, topsep=0pt,parsep=0pt,partopsep=0pt,noitemsep

### v1

@@ Abstract
Training deep neural networks (DNNs) is costly. Fortunately, Nvidia Ampere and Hopper GPUs can accelerate matrix multiplications twice as fast as a dense equivalent by implementing 2:4 sparsity. However, previous STE-based 2:4 pre-training methods (e.g. STE with hard-thresholding, SR-STE) suffer from optimization difficulties because of discontinuous pruning function. In this study, we comprehensively analyse the bottleneck of traditional N:M sparse training and recognize three drawbacks with discontinuity: incorrect descending direction, inability to predict the amount of descent and sparse mask oscillation. In the light of this statement, we propose S-STE, a simple yet powerful 2:4 training method that contains two parts: to continuously project weights to be 2:4 sparse, and to rescale sparse weights with a per-tensor fixed scaling factor. Besides, we adopt minimum-variance unbiased estimation for activation gradient and FP8 quantization for whole process. Results show that our method surpasses previous 2:4 pre-training recipes and is comparable even with full parameter models.
@@ 1 Introduction
Large scale transformers have achieved many impressive results such as chatbots Touvron et al. (2023), text-to-video generation Liu et al. (2024), and robot manipulation Brohan et al. (2023). However, the pre-training of these models is extremely expensive, typically requiring thousands of GPUs to train for months Brown et al. (2020). One possible way to accelerate deep learning computation is sparsity. N:M sparsity Mishra et al. (2021) is a hardware-friendly sparsity pattern, where every group of $M$ dimensions only has $N$ non-zero entries. Nvidia Ampere GPUs can multiply a 2:4 sparse matrix with a dense matrix, twice as fast as multiplying two dense matrices.
@@ 3.1 Phenomenon 1: incorrect descending direction
$F(\boldsymbol{\mathbf{w}}_{k})-F(\boldsymbol{\mathbf{w}}_{k+1})\approx(\nabla_{\boldsymbol{\mathbf{w}}_{k}}F(\boldsymbol{\mathbf{w}}_{k}))^{\top}(\boldsymbol{\mathbf{w}}_{k}-\boldsymbol{\mathbf{w}}_{k+1})=\alpha_{k}\left\lVert\boldsymbol{\mathbf{w}}_{k}-\boldsymbol{\mathbf{w}}_{k+1}\right\rVert^{2}\geq 0.$ (4)
@@ 4 Methodology
$\gamma$ validation loss test BLEU
0 4.007 26.3
@@
In this section we propose a training algorithm (smooth straight-through estimator, S-STE) that contains two main parts combined with STE: 2:4 specific soft-thresholding, and fixed weight rescaling . They together work as the sparsifying function described in Sec. 2: $\tilde{\boldsymbol{\mathbf{w}}}=S(\boldsymbol{\mathbf{w}})=\beta S_{soft}(\boldsymbol{\mathbf{w}})$. Results in Fig. 2(c)(d) and 4(d) show that S-STE successfully overcome the three curses of discontinuity.
@@ 4.2 Fixed weight rescaling $\beta$
S-STE $\operatorname{MVUE}$ $(S(\mathbf{W})^{\top})$ $\operatorname{MVUE}$ $(\nabla_{\mathbf{Z}}^{\top})$ loss
3.3948
3.4333
3.4644
3.4773
3.448
SR-STE 3.4739
@@
$\beta$ recipe test bleu validation loss average epoch loss
no scaling 25.28 4.044 4.67
@@
Figure 5: Standard deviation of weight gradients and input activation gradients of different layers of GPT-2 large 774M model. (a)(b) shows the result of sparsifying both $S(\boldsymbol{\mathbf{W}}_{l})^{\top}$ and $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}^{\top}$. (c) shows the ratio of standard deviation of dense and sparse model in (a) and (b) respectively. (d) shows the result of only sparsifying $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}^{\top}$. Note that $\sigma(\nabla_{\mathbf{X}_{l}})$ of dense model and sparsifying $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}^{\top}$ model are the same here, which we do not depict.
@@ 5.1 Minimum-variance unbiased estimation
To accelerate the backward propagation, Chmiel et al. (2023) suggest using a minimum-variance unbiased estimator (MVUE). For every linear layer $\boldsymbol{\mathbf{Z}}_{l}=\boldsymbol{\mathbf{X}}_{l}S(\boldsymbol{\mathbf{W}}_{l})^{\top}$, there are two matrix multiplications of the backward pass in total: $\nabla_{\boldsymbol{\mathbf{X}}_{l}}=\nabla_{\boldsymbol{\mathbf{Z}}_{l}}S(\boldsymbol{\mathbf{W}}_{l})$ and $\nabla_{\boldsymbol{\mathbf{W}}_{l}}=\nabla_{\boldsymbol{\mathbf{Z}}_{l}}^{\top}\boldsymbol{\mathbf{X}}_{l}$, where $\boldsymbol{\mathbf{X}}_{l}$ is the input of the $l$-th layer, $\boldsymbol{\mathbf{W}}_{l}$ and $\boldsymbol{\mathbf{Z}}_{l}$ are the weight matrix and output activation. We conduct MVUE on both two matrix multiplications and compare their results: $\nabla_{\boldsymbol{\mathbf{X}}_{l}}=\nabla_{\boldsymbol{\mathbf{Z}}_{l}}\operatorname{MVUE}(S(\boldsymbol{\mathbf{W}}_{l})^{\top})^{\top}$ and $\nabla_{\boldsymbol{\mathbf{W}}_{l}}=\operatorname{MVUE}(\nabla_{\boldsymbol{\mathbf{Z}}_{l}}^{\top})\boldsymbol{\mathbf{X}}_{l}$. Specifically, we choose $S(\boldsymbol{\mathbf{W}}_{l})$ and $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}$ because they both have built-in sparsity Lin et al. (2020). However, we only choose to sparsify the latter one. Firstly, it is proven by Hu et al. (2024) and Chmiel et al. (2023) that minimum loss of accuracy is guaranteed for MVUE on $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}$. Secondly, using MVUE on $S(\boldsymbol{\mathbf{W}}_{l})$ will make errors accumulate along the back propagation, and results in large standard deviation for the first few layers; see Fig. 5. Besides, results in Table 2 also show minimum loss of accuracy of MVUE on $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}$ while obvious accuracy loss on $S(\boldsymbol{\mathbf{W}}_{l})$. Thus, we choose to sparsify only $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}$ in the backward pass.
@@ 5.2 FP8 training
6 Experiments
For Transformer, we train Transformer-base models on WMT 14 En-De dataset Bojar et al. (2014) with fairseq Ott et al. (2019) codebase and evaluate it with BLEU Papineni et al. (2002) scores. For DeiT, we pre-train Deit-small model for ImageNet-1K Deng et al. (2009) classification task. For GPT-2, we pre-train GPT-2 124M, 350M and 774M models on OpenWebText Gokaslan and Cohen (2019) and evaluate it on GLUE Wang et al. (2019) and SQuAD Rajpurkar et al. (2016) benchmarks. We also compare our method with state-of-the-art 2:4 training methods (SR-STE Zhou et al. (2021), Bi-Mask Zhang et al. (2023) and STEP Lu et al. (2023)). All experiments are done on RTX 4090 machines.
Table 5: Comparison between GLUE scores of different pre-train methods on GPT-2 models. We use full parameters to evaluate dense and SR-STE+DF models, and half parameters (2:4 sparse) to evaluate S-STE models.
params method val loss avg score
124M Dense 2.907 $73.9\pm 1.1$
SR-STE $3.076$ $72.6\pm 0.2$
S-STE $2.984$ $74.1\pm 0.4$
350M Dense 2.618 $76.3\pm 0.1$
SR-STE $2.718$ $76.3\pm 0.4$
S-STE $2.713$ $76.9\pm 0.6$
774M Dense 2.493 $76.2\pm 0.4$
SR-STE+DF $2.564$ $77.1\pm 0.4$
S-STE $2.547$ $78.6\pm 0.8$
Machine translation
@@ 6 Experiments
We further investigate the effectiveness of S-STE to train DeiT-small on ImageNet-1k; see Table 6. Results show S-STE also achieve the best performance among different methods, with only has 1.4% degradation from the dense model. Notably, S-STE surpasses SOTA 2:4 training method Bi-Mask on this task (0.9% top1 accuracy improvement) and popular SR-STE method (2.8% top1 accuracy improvement).
@@
We compare S-STE with dense, normal SR-STE and SR-STE with dense fine-tuning Hu et al. (2024) (SR-STE+DF) models on GLUE and SQuAD tasks. The SR-STE+DF models first use SR-STE to train a 2:4 sparse model, and switch to dense training for the last 1/6 iters of pre-training (which stands for “dense fine-tune”). In downstream tasks it also use dense parameters to make predictions, similar to dense models. Results in Table 5, 8 and 9 show that S-STE completely surpasses SR-STE on both tasks. Even for SR-STE+DF models, S-STE still have an advantage, with an improvement of 1.5 on GLUE average and 1.2/0.9 on SQuAD for GPT-2 774M.
method test acc1 test acc5 test loss
dense 79.9 95 0.887
SR-STE Zhang et al. (2023) 75.7 - -
Bi-Mask Zhang et al. (2023) 77.6 - -
S-STE $78.5$ $94.4$ $0.938$
@@
We explore the effectiveness of S-STE, MVUE, and FP8 separately in this part. We pre-train DeiT-small model on ImageNet-1K dataset for image classification. Combinations of these partitions in Table 7 show that: 1) FP8 training has little affect on pre-training accuracy (0.2% of acc1); 2) MVUE leads to minimal loss of performance (0.1% of acc1).
Table 7: Experimental result of S-STE, MVUE and FP8 training on DeiT-small.
S-STE $\operatorname{MVUE}(\nabla_{\boldsymbol{\mathbf{Z}}}^{\top})$ FP8 test acc1 test acc5
79.9 95
79.7 94.9
@@
Table 8: SQuAD scores of different sizes and pre-training methods on GPT-2. Similar to Table 5, we use 2:4 sparse weights to evaluate S-STE model, while dense parameters to evaluate the rest.
params method EM F1
124M Dense 67.6 78.8
SR-STE+DF 67.5 78.5
S-STE $68$ $78.8$
350M Dense 73.2 83.6
SR-STE+DF 71.9 82.4
S-STE $72.7$ $82.7$
774M Dense 74.3 84.9
SR-STE+DF 74.3 84.6
S-STE $75.5$ $85.5$
@@ 7 Related work
While 16-bit float tensors are widely used in pre-training, FP8 – where float numbers stored in 8 bits – is a popular quantization methods which theoretically accelerates GEMMs up to 4x faster than its fp32 counterparts and 2x faster than its FP16/BF16 counterparts Micikevicius et al. (2022); Perez et al. (2023); van Baalen et al. (2023); Wang et al. (2018); Noune et al. (2022). With e3m4 data format used in forward and e5m2 format Sun et al. (2019) in backward, pre-trained models can achieve minimum loss of accuracy while greatly boosting the efficiency of training.
8 Conclusion
In this study we discuss the importance of pruning continuity in effective 2:4 sparse pre-training. We analyse the drawback of traditional hard-thresholding pruning function and its variation (SR-STE) and argue that the main limits being discontinuity. Based on our analysis and soft-thresholding for channel pruning, we propose S-STE method, which prunes weights in a continuous manner. Experiments show that our method surpasses previous state-of-the-art methods on a wide range of tasks.

### v3

@@ Abstract
Training deep neural networks (DNNs) is costly. Fortunately, Nvidia Ampere and Hopper GPUs can accelerate matrix multiplications twice as fast as a dense equivalent by implementing 2:4 sparsity. However, previous STE-based 2:4 pre-training methods (e.g. STE with hard-thresholding, SR-STE) suffer from optimization difficulties because of discontinuous pruning function. In this study, we comprehensively analyse the bottleneck of traditional N:M sparse training and recognize three drawbacks with discontinuity: incorrect descending direction, inability to predict the amount of descent and sparse mask oscillation. In light of this, we propose S-STE, a simple yet powerful 2:4 training method that contains two parts: to continuously project weights to be 2:4 sparse, and to rescale sparse weights with a per-tensor fixed scaling factor. Besides, we adopt minimum-variance unbiased estimation for activation gradient and FP8 quantization for whole process. Results show that our method surpasses previous 2:4 pre-training recipes and is comparable even with full parameter models. Our toolkit is available at https://github.com/huyz2023/2by4-pretrain.
@@ 1 Introduction
Large scale transformers have achieved many impressive results such as chatbots Touvron et al. (2023), text-to-video generation Liu et al. (2024), and robot manipulation Zitkovich et al. (2023). However, the pre-training of these models is extremely expensive, typically requiring thousands of GPUs to train for months Brown et al. (2020). One possible way to accelerate deep learning computation is sparsity. N:M sparsity Mishra et al. (2021) is a hardware-friendly sparsity pattern, where every group of $M$ dimensions only has $N$ non-zero entries. Nvidia Ampere GPUs can multiply a 2:4 sparse matrix with a dense matrix, twice as fast as multiplying two dense matrices.
@@ 3.1 Phenomenon 1: incorrect descending direction
$F(\boldsymbol{\mathbf{w}}_{k})-F(\boldsymbol{\mathbf{w}}_{k+1})\approx(\nabla_{\boldsymbol{\mathbf{w}}_{k}}F(\boldsymbol{\mathbf{w}}_{k}))^{\top}(\boldsymbol{\mathbf{w}}_{k}-\boldsymbol{\mathbf{w}}_{k+1})=\alpha_{k}\left\lVert\nabla_{\boldsymbol{\mathbf{w}}_{k}}F(\boldsymbol{\mathbf{w}}_{k})\right\rVert^{2}\geq 0.$ (4)
@@ 4 Methodology
$\gamma$ Val loss Test BLEU
0 4.007 26.30
@@
In this section we propose a training algorithm (smooth straight-through estimator, S-STE) that contains two main parts combined with STE: 2:4 specific soft-thresholding, and fixed weight rescaling . They together work as the sparsifying function described in Sec. 2: $\tilde{\boldsymbol{\mathbf{w}}}=S(\boldsymbol{\mathbf{w}})=\beta S_{soft}(\boldsymbol{\mathbf{w}})$. Results in Fig. 2(c)(d) and 4(d) show that S-STE successfully overcome the three curses of discontinuity. Notably, flip rate curves of S-STE are surprisingly consistent with their dense counterparts, indicating that S-STE is more natural and feasible than SR-STE.
@@ 4.2 Fixed weight rescaling $\beta$
$\beta$ Recipe Test BLEU Val loss Avg epoch loss
No scaling 25.28 4.044 4.670
@@ 5.1 Minimum-variance unbiased estimation
To accelerate the backward propagation, Chmiel et al. (2023) suggest using a minimum-variance unbiased estimator (MVUE). For every linear layer $\boldsymbol{\mathbf{Z}}_{l}=\boldsymbol{\mathbf{X}}_{l}S(\boldsymbol{\mathbf{W}}_{l})^{\top}$, there are two matrix multiplications of the backward pass in total: $\nabla_{\boldsymbol{\mathbf{X}}_{l}}=\nabla_{\boldsymbol{\mathbf{Z}}_{l}}S(\boldsymbol{\mathbf{W}}_{l})$ and $\nabla_{\boldsymbol{\mathbf{W}}_{l}}=\nabla_{\boldsymbol{\mathbf{Z}}_{l}}^{\top}\boldsymbol{\mathbf{X}}_{l}$, where $\boldsymbol{\mathbf{X}}_{l}$ is the input of the $l$-th layer, $\boldsymbol{\mathbf{W}}_{l}$ and $\boldsymbol{\mathbf{Z}}_{l}$ are the weight matrix and output activation. We conduct MVUE on both two matrix multiplications and compare their results: $\nabla_{\boldsymbol{\mathbf{X}}_{l}}=\nabla_{\boldsymbol{\mathbf{Z}}_{l}}\operatorname{MVUE}(S(\boldsymbol{\mathbf{W}}_{l})^{\top})^{\top}$ and $\nabla_{\boldsymbol{\mathbf{W}}_{l}}=\operatorname{MVUE}(\nabla_{\boldsymbol{\mathbf{Z}}_{l}}^{\top})\boldsymbol{\mathbf{X}}_{l}$. Specifically, we choose $S(\boldsymbol{\mathbf{W}}_{l})$ and $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}$ because they both have built-in sparsity Lin et al. (2020). However, we only choose to sparsify the latter one. Firstly, it is proven by Hu et al. (2024) and Chmiel et al. (2023) that minimum loss of accuracy is guaranteed for MVUE on $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}$. Secondly, using MVUE on $S(\boldsymbol{\mathbf{W}}_{l})$ will make errors accumulate along the back propagation, and results in large standard deviation of gradient for the first few layers. Besides, results in Table 3 also show minimum loss of accuracy of MVUE on $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}$ while obvious accuracy loss on $S(\boldsymbol{\mathbf{W}}_{l})$. Thus, we choose to sparsify only $\nabla_{\boldsymbol{\mathbf{Z}}_{l}}$ in the backward pass.
S-STE $\operatorname{MVUE}(S(\mathbf{W})^{\top})$ $\operatorname{MVUE}(\nabla_{\mathbf{Z}}^{\top})$ comment loss
- dense 3.3948
- SR-STE 3.4739
3.4333
3.4644
3.4773
3.4480
@@ 5.2 FP8 training
6 Experiments
For Transformer, we train Transformer-base models on WMT 14 En-De dataset Bojar et al. (2014) with fairseq Ott et al. (2019) codebase and evaluate it with BLEU Papineni et al. (2002) scores. For DeiT, we pre-train Deit-small model for ImageNet-1K Deng et al. (2009) classification task. For GPT-2, we pre-train GPT-2 124M, 350M and 774M models on OpenWebText Gokaslan and Cohen (2019) and evaluate it on GLUE Wang et al. (2018a) and SQuAD Rajpurkar et al. (2016) benchmarks. We also compare our method with state-of-the-art 2:4 training methods (SR-STE Zhou et al. (2021), Bi-Mask Zhang et al. (2023) and STEP Lu et al. (2023)). The pre-training and evaluation scripts are publicly available at https://github.com/thu-ml/2by4-pretrain-acc-examples.
Machine translation
@@ 6 Experiments
Model Method Test acc1 Test acc5
DeiT-tiny Dense 72.2 91.1
SR-STE Zhang et al. (2023) 67.8 88.6
S-STE $68.5$ $88.9$
DeiT-small Dense 79.9 95
SR-STE Zhang et al. (2023) 75.7 -
Bi-Mask Zhang et al. (2023) 77.6 -
S-STE $78.5$ $94.4$
@@
We further investigate the effectiveness of S-STE to train DeiT-tiny and DeiT-small on ImageNet-1k; see Table 5. Results show S-STE also achieve the best performance among different methods, with only has 1.4% degradation from the dense model. Notably, S-STE surpasses SOTA 2:4 training method Bi-Mask on this task (0.9% top1 accuracy improvement) and popular SR-STE method (2.8% top1 accuracy improvement).
@@
We compare S-STE with dense, normal SR-STE and SR-STE with dense fine-tuning Hu et al. (2024) (SR-STE+DF) models on GLUE and SQuAD tasks. The SR-STE+DF models first use SR-STE to train a 2:4 sparse model, and switch to dense training for the last 1/6 iters of pre-training (which stands for “dense fine-tune”). In downstream tasks it also use dense parameters to make predictions, similar to dense models. Results in Table 6 and 9 show that S-STE completely surpasses SR-STE on both tasks. Even for SR-STE+DF models, S-STE still have an advantage, with an improvement of 1.5 on GLUE average and 1.2/0.9 on SQuAD for GPT-2 774M.
Table 6: SQuAD and GLUE scores of different sizes and pre-training methods on GPT-2. We use 2:4 sparse weights to evaluate S-STE model, while dense parameters to evaluate the rest. Of note, SR-STE denotes the original SR-STE workflow (without backward MVUE), and “T-SR-STE+DF” denotes the combination of transposable SR-STE & backward MVUE & sparse-dense training workflow, proposed by Hu et al. (2024). S-STE settings here include backward MVUE & FP8 training.
Params Pre-train Fine-tune Pre-train val loss SQuAD GLUE@Avg
EM F1
124M Dense Dense 2.907 67.6 78.8 $73.9\pm 1.1$
T-SR-STE+DF Hu et al. (2024) Dense 2.952 67.5 78.5 $74.3\pm 0.5$
T-SR-STE Dense 3.076 66.3 77.2 $72.6\pm 0.2$
SR-STE Dense 2.982 66.2 77.5 $73.8\pm 0.3$
S-STE S-STE 2.984 68 78.8 $74.1\pm 0.4$
350M Dense Dense 2.618 73.2 83.6 $76.3\pm 0.1$
T-SR-STE+DF Hu et al. (2024) Dense 2.688 71.9 82.4 $77.1\pm 0.2$
T-SR-STE Dense 2.718 72.3 82.6 $76.3\pm 0.4$
SR-STE Dense 2.690 72.0 82.4 $76.8\pm 0.4$
S-STE S-STE 2.713 72.2 82.7 $76.9\pm 0.6$
774M Dense Dense 2.493 74.3 84.9 $76.2\pm 0.4$
T-SR-STE+DF Hu et al. (2024) Dense 2.564 74.3 84.6 $77.1\pm 0.4$
S-STE S-STE 2.547 75.5 85.5 $78.6\pm 0.8$
Table 7: Different fine-tuning results on GLUE and SQuAD.
Model Downstream task Pre-train Fine-tune Avg score
GPT-2 124M GLUE S-STE Hard-thresholding $73.9\pm 0.6$
GLUE S-STE S-STE $74.1\pm 0.4$
SQuAD S-STE Hard-thresholding $67.6/78.6$
SQuAD S-STE S-STE $68/78.8$
Fine-tuning
We illustrate the viability of S-STE for fine-tuning a pre-trained model, presenting a coherent workflow of accelerating both training and inference (dense fine-tuning cannot produce a sparse model for inference acceleration); see Table 7, 6.
@@
In this part, We explore the effectiveness of S-STE, MVUE, and FP8 separately. We pre-train DeiT-small model on ImageNet-1K dataset for image classification. Combinations of these partitions in Table 8 show that: 1) FP8 training has little affect on pre-training accuracy (0.2% of acc1); 2) MVUE leads to minimal loss of performance (0.1% of acc1).
Table 8: Experimental result of S-STE (soft-thresholding and weight rescaling), MVUE and FP8 training with DeiT-small on ImageNet-1K.
Soft- thresholding Weight rescaling $\operatorname{MVUE}(\nabla_{\mathbf{Z}}^{\top})$ FP8 Comment Test acc1 Test acc5
- - Dense 79.9 95
- - Dense; FP8 79.7 94.9
Hard-thresholding 77.7 93.9
@@
Acceleration
For acceleration, we measure the acceleration ratio of a typical GPT-2 model using implementation from Hu et al. (2024). Note that on H100 GPUs, FP8 2:4-spMM kernel turns out to be unsatisfying; see Appendix A.4. Consequently, we fall back to use RTX3090 GPUs with FP16 training. For inference, we achieve 1.53x speedup with FFN layer and 1.23x speedup with the network; for pre-training, we achieve 1.32x speedup for FFN layer and 1.18x speedup for the network (Appendix A.3).
@@ 7 Related work
While 16-bit float tensors are widely used in pre-training, FP8 – where float numbers stored in 8 bits – is a popular quantization methods which theoretically accelerates GEMMs up to 4x faster than its fp32 counterparts and 2x faster than its FP16/BF16 counterparts Micikevicius et al. (2022); Perez et al. (2023); van Baalen et al. (2023); Wang et al. (2018b); Noune et al. (2022). With e3m4 data format used in forward and e5m2 format Sun et al. (2019) in backward, pre-trained models can achieve minimum loss of accuracy while greatly boosting the efficiency of training.
8 Conclusions and future work
In this study we discuss the importance of pruning continuity in effective 2:4 sparse pre-training. We analyse the drawback of traditional hard-thresholding pruning function and its variation (SR-STE) and argue that the main limits being discontinuity. Based on our analysis and soft-thresholding for channel pruning, we propose S-STE, which prunes weights in a continuous manner. Experiments show that our method surpasses previous state-of-the-art methods on a wide range of tasks.
Our proposed S-STE approach primarily targets linear layers within FFN networks. Nevertheless, QKV projection layers necessitate further exploration to devise an effective dynamic sparse training strategy that harmonizes with attention mechanisms. Furthermore, our current choice of continuous pruning function represents only one possible solution; alternative, smoother pruning functions may be necessary to achieve improved continuity and mitigate potential discontinuities.